The time it takes to do homework includes a fixed amount of time to prepare plus a constant amount of time per problem. If a student can do 5 homework problems in 40 minutes, and 10 problems in 70 minutes, how many minutes will 25 problems take?
Answer by mananth(16946)

The time it takes to do homework includes a fixed amount of time to prepare plus a constant amount of time per problem. If a student can do 5 homework problems in 40 minutes, and 10 problems in 70 minutes, how many minutes will 25 problems take?
Let the fixed time taken be x
The time to solve a problem be y
x+5y=40------------1
x+10y=70-----------2
subtract equation 2 from 1
(x+5y) - (x+10y) = 40-70
x+5y -x-10y = -30
-5y=-30
y=-30/-5
y=4. Time taken to solve each problem is 4 minutes
x+5y=40 substitute the value of y in the equation
x+20=40
x=40-20 = 20 minutes will be the fixed time
25 problems will take (20 minutes + 4 *25)
= 120 minutes or 2 hours